# Real-Time Position-Aware View Synthesis from Single-View Input

### v1

@@ Abstract.
Recent advancements in view synthesis have significantly enhanced immersive experiences across various computer graphics and multimedia applications, including telepresence, and entertainment. By enabling the generation of new perspectives from a single input view, view synthesis allows users to better perceive and interact with their environment. However, many state-of-the-art methods, while achieving high visual quality, face limitations in real-time performance, which makes them less suitable for live applications where low latency is critical. In this paper, we present a lightweight, position-aware network designed for real-time view synthesis from a single input image and a target camera pose. The proposed framework consists of a Position Aware Embedding, modeled with a multi-layer perceptron, which efficiently maps positional information from the target pose to generate high dimensional feature maps. These feature maps, along with the input image, are fed into a Rendering Network that merges features from dual encoder branches to resolve both high level semantics and low level details, producing a realistic new view of the scene. Experimental results demonstrate that our method achieves superior efficiency and visual quality compared to existing approaches, particularly in handling complex translational movements without explicit geometric operations like warping. This work marks a step toward enabling real-time view synthesis from a single image for live and interactive applications.
@@ 1. Introduction
Telepresence systems (Dima and Sjöström, 2021) enable users to interact with remote environments in real time, offering a wide range of applications, particularly in sectors such as heavy machinery, where remote operation from a safe distance is crucial. These systems have the potential to facilitate seamless natural interaction with remote environments. Current telepresence technologies typically rely on head-mounted displays (HMDs), high-speed networks, and consumer or industrial-grade cameras to deliver live video feeds of the environment. While 2D video feeds provide some pictorial information, the lack of depth perception significantly limits situational awareness, which is vital for effective remote operation.
The challenge to have an understanding of a remote site’s structure may be address by enabling motion parallax for the user. I.e., the user moves the head forth and back to get a notion of the scene depth. This requires access to views from different perspectives, where camera positions and poses need to be continuous for a smooth transition and a high quality experience. Therefore, it is imperative to render continuous perspective views accurately and in real time for applications such as augmented reality (AR), virtual reality (VR), and telepresence systems.
@@
The synthesis of views from perspectives continuous in camera position and pose requires sufficient information of the scene. Early methods employed view interpolations and depth-image-based rendering, but both these require multiple capture devices resulting in calibration issues and high communication capacity requirements. Therefore, generating new views from a single input image and a specified camera pose has emerged as a key task in rendering. A variety of approaches, such as Neural Radiance Fields (NeRF) (Mildenhall et al., 2021) and more recent transformer-based architectures like NViST (Jang and Agapito, 2024), have been proposed to tackle this problem. These methods leverage both synthetic and real-world datasets to produce high-quality novel views. However, most state-of-the-art methods either require multi-view inputs or are computationally intensive, making them unsuitable for real-time applications where low latency is critical.
In this work, we build upon the growing body of research on single-image-based view synthesis. We introduce a learning based position embedding method combined with a rendering network that efficiently synthesizes high-quality views of arbitrary camera location and pose using fewer computational resources. Our method maintains high fidelity even when handling large translational shifts, making it ideal for real-time applications and practical deployment on consumer devices. To train and test our method we also create our own synthetic dataset using Blender.
Unlike other methods, such as NViST, which employs transformer architectures, or SE3D (Koh et al., 2023), which utilizes large-scale generative-adversarial networks (GANs) to address occlusions and complex scenes, our network is designed to be lightweight and computationally efficient. Despite its smaller footprint, it excels in handling scenarios where translational movements dominate, enabling real-time inference. In summary, our main contributions are as follows:
A lightweight network with a positional embedding module that can generate novel views within a given baseline without requiring explicit operations.
@@
A positional encoding scheme similar to NeRF, which is combined with an multi-layer perceptron (MLP), to create a more generalizable framework capable of handling different scenes.
Our method achieves real-time performance, with inference times of approximately 8.6 milliseconds, corresponding to over 100 frames per second (FPS) for images at a resolution of 512x512 pixels.
@@
The remainder of this paper is organized as follows: Section 2 reviews state-of-the-art methods aimed at view synthesis. Section 3, and 4 describes the proposed approach for small baseline and large baseline rendering. Section 5 provides a comparative analysis of the results, including quality and speed evaluations. Finally, conclusions are summarized in Section 7.
@@ 2. Background
View synthesis has recently become a pivotal research area in immersive telepresence, aiming to generate unseen perspectives of a remote scene from a limited set of input images. Traditional approaches, such as structure-from-motion (SfM) and multi-view stereo (MVS), rely on geometric consistency between input views to reconstruct novel viewpoints (Seitz and Dyer, 1999). While these methods have shown effectiveness in controlled environments, their performance deteriorates in complex real-world scenarios and often necessitates multiple input views to achieve high-quality results. In light of these limitations, more recent methods focus on synthesizing novel views from a sparse or single image. This section outlines these advancements and their respective contributions to the field of view synthesis.
2.1. Scene Regression Approaches
@@
Recent advancements in view synthesis have centered around methods that aim to represent a scene from sparse image inputs, particularly those leveraging neural implicit representations such as NeRF (Mildenhall et al., 2021). A NeRF encodes a scene as a continuous volumetric function, allowing novel view generation through volumetric rendering. Despite its capacity to produce highly photorealistic images, NeRF suffers from computational inefficiencies, particularly in terms of inference time. Several techniques have been proposed to mitigate these limitations, such as hash encoding in InstantNGP (Müller et al., 2022) and multi-scale representations in Mip-NeRF (Barron et al., 2021). While these improvements accelerate rendering, they retain significant reliance on MLPs, which remain computationally expensive.
@@
To address these concerns, methods like Plenoxels (Fridovich-Keil et al., 2022) have emerged, leveraging spherical harmonics to represent a scene without the need for deep neural networks. Though Plenoxels achieves faster optimization than many NeRF-based methods, it still falls short in terms of memory efficiency. A more recent innovation, Gaussian Splatting (Kerbl et al., 2023), improves memory and rendering efficiency by employing high-dimensional Gaussian kernels to represent the scene. However, its optimal performance still depends on the availability of multiple input views.
2.2. Light Field Reconstruction and Rendering
@@
Light field (LF) representation encodes both the spatial and angular information of a scene, often described by an HxWxNxN structure, where HxW refers to the spatial resolution and NxN denotes the angular resolution (Levoy and Hanrahan, 1996). This representation facilitates tasks like viewpoint manipulation (Zhou et al., 2020). LF reconstruction methods aim to generate a dense NxN light field from a single input view or a few input views, eliminating the need for complex multi-camera setups.
A common LF reconstruction approach utilizes convolutional neural networks (CNNs) to estimate scene depth (Srinivasan et al., 2017), followed by warping operators to generate target views. To address warping artifacts, subsequent refinement networks have been developed (Zhou et al., 2020, 2021). Other methods exploit depth estimation or optical flow (Cun et al., 2019), appearance flow (Ivan et al., 2019), or a combination of both (Bae et al., 2021) to synthesize the light field.
In more recent developments, GANs (Chen et al., 2020; Chandramouli et al., 2020) and auto-encoders (Han and Xiang, 2022) have demonstrated effectiveness in reconstructing dense light fields from a single image input. Additionally, LFSphereNet (Gond et al., 2023) extends LF reconstruction to spherical light fields using a single omnidirectional image. Furthermore, to generate the novel views from the LFs, the algorithms like dynamically reparameterized LFs (Isaksen et al., 2000) and per-view disk-based blending (Overbeck et al., 2018) have to be utilized which have higher memory footprint, making the whole process unsuitable for dynamic environments and low memory systems.
Our proposed method, PLFNet (described in Sec 3), contributes to LF reconstruction and rendering by directly synthesizing views within a local LF volume, thus eliminating the need of additional rendering algorithms.
2.3. Learning-Based Approaches
Learning-based methods seek to overcome the limitations of both scene regression and light field rendering approaches, which include constraints such as low baselines, sparse image input requirements, and slow rendering times. Several models, including PixelNeRF (Yu et al., 2021) and VisionNeRF (Lin et al., 2023), have been developed to generalize NeRF to novel scenes by conditioning the network on image features. Despite their promise, these methods often face challenges such as scale ambiguity and difficulty in handling large translational movements.
@@ 2.3. Learning-Based Approaches
To address these challenges, the diffusion model based approach in 3DiM (Watson et al., 2022) employs pose guided diffusion to improve performance, particularly in scenarios with large translational movement. Fine-tuning diffusion models on large-scale synthetic datasets (Deitke et al., 2023) has also shown potential (Liu et al., 2023, 2024), although these methods are hindered by slow inference times.
@@
An alternative strategy involves using multiplane image (MPI) representations, which approximate a scene using a series of parallel planes at fixed depths, each containing an image and alpha map. This representation allows novel view generation based on a fixed (Zhou et al., 2018; Flynn et al., 2019; Tucker and Snavely, 2020; Li and Kalantari, 2020) or variable (Li and Kalantari, 2020) number of planes. While MPI-based methods achieve superior visual quality, their computational demands remain high, especially during MPI layer prediction. Large-scale GANs (Koh et al., 2023) have also been explored for their ability to render fast and handle large baselines, though they introduce significant memory overhead, limiting their practicality.
More recent innovations, such as NViST (Jang and Agapito, 2024), incorporate vision transformers (ViTs) (Dosovitskiy, 2020) into view synthesis architectures, demonstrating strong generalization capabilities in real-world settings. By leveraging masked autoencoders (MAEs) and adaptive layer normalization conditioned on camera parameters, NViST delivers high-quality results, even for out-of-distribution scenes.
Our proposed method, PLFNet+ (described in Sec 4), contributes to the learning-based category by extending the views from a single input image while diverging from diffusion and transformer-based methods. By utilizing a lightweight design, PLFNet+ reduces computational overhead, enabling real-time rendering. This efficiency makes it particularly suited for dynamic environments where fast processing is crucial, addressing key challenges in the field of view synthesis.
3. Method: Light Field Reconstruction - PLFNet
We first define a Position Aware Light Field Reconstruction network called ”PLFNet” which aims to reconstruct the LF from a single input image by iterative reconstructing individual sub-aperture images. The source image serves as the basis for reconstructing the surrounding LF , where x represents the angular coordinate , and signifies the center of the LF. Each sub-aperture image at location , specified as , undergoes reconstruction via:
This two stage pipeline has been shown in Figure 1

### v2

@@ Abstract.
Recent advancements in view synthesis have significantly enhanced immersive experiences across various computer graphics and multimedia applications, including telepresence and entertainment. By enabling the generation of new perspectives from a single input view, view synthesis allows users to better perceive and interact with their environment. However, many state-of-the-art methods, while achieving high visual quality, face limitations in real-time performance, which makes them less suitable for live applications where low latency is critical. In this paper, we present a lightweight, position-aware network designed for real-time view synthesis from a single input image and a target camera pose. The proposed framework consists of a Position Aware Embedding, which efficiently maps positional information from the target pose to generate high dimensional feature maps. These feature maps, along with the input image, are fed into a Rendering Network that merges features from dual encoder branches to resolve both high and low level details, producing a realistic new view of the scene. Experimental results demonstrate that our method achieves superior efficiency and visual quality compared to existing approaches, particularly in handling complex translational movements without explicit geometric operations like warping. This work marks a step toward enabling real-time live and interactive telepresence applications.
@@ 1. Introduction
Telepresence systems (Dima and Sjöström, 2021; Chandan et al., 2021; Sakashita et al., 2023), allow users to interact with remote environments in real-time, offering a wide range of applications, particularly in sectors such as heavy machinery, where remote operation from a safe distance is crucial. These systems have the potential to facilitate seamless natural interaction with remote environments. Current real-time telepresence technologies (Chandan et al., 2021; Sakashita et al., 2023) typically rely on head-mounted displays (HMDs), high-speed networks, and consumer or industrial-grade cameras to deliver live video feeds of the environment. While live 2D video feeds provide visual information, they fall short in offering depth perception (Chandan et al., 2021). The lack of depth perception significantly limits situational awareness (Dima et al., 2020), which is vital for effective remote operation.
The lack of depth perception that a 2D video feed provides can be leveraged by binocular parallax and motion parallax as depth cues. Recently, SpinVR (Konrad et al., 2017) utilized a stereoscopic virtual reality setup to provide binocular parallax to the users. However, this still lacks motion parallax. When a user moves their viewpoint in space, the relative motion of objects within the scene provides the brain with cues about the depth. Thus, rendering new views at different positions can simulate motion parallax, improving depth perception and enhancing situational awareness in remote operations. For effective depth perception and smooth interaction in real-time, synthesizing new views, while maintaining low latency, becomes essential as it enables both binocular and motion parallaxes.
Synthesizing views from different perspectives requires sufficient information of the scene. While 360-degree cameras are used in some telepresence system (Chandan et al., 2021), synthesizing new views from these omnidirectional images (ODIs) as done in works like PanoSynthVR (Waidhofer et al., 2022) introduces additional complexities. The distortion inherent in ODIs and the required high bandwidth for transferring such data make these methods less practical for real-time use. Instead, focusing on planar images captured from the viewpoint of the user offers a more efficient solution.
We consider a telepresence application as the primary use case for our view synthesis method, illustrated in Figure 1. In this setting, an omnidirectional image (ODI) is captured from remote scene. A viewport image is extracted from the ODI based on the current gaze direction of the user, while the relative head movement (i.e., translation) is used to define a target position. The extracted image, target position, and target rotation are then passed to our network, which synthesizes a new view at the intended offset. The resulting image is rendered as the updated viewport, enabling motion parallax without requiring a full 3D reconstruction or stitching it back into the ODI. This patch based rendering pipeline eliminates the distortions which are present in ODIs, reduces transmission bandwidth, and allows for efficient local view synthesis.
By synthesizing views directly from these extracted planar patches, we avoid the distortions inherent in ODIs and reduce the bandwidth needed for transmitting high resolution spherical imagery. This patch based synthesis pipeline supports real-time performance and enables localized motion parallax around the viewpoint of the user. However, such ODI-based workflows inherently embed orientation within the extracted view direction, requiring only translational offsets for rendering. To extend support to general planar RGB inputs such as monocular cameras not embedded in a spherical system, our method also incorporates full 6DoF inputs, including yaw, pitch, and roll. This makes our framework broadly applicable beyond the illustrated telepresence scenario.
@@
Early view synthesis methods (Ahn and Kim, 2013; Yao et al., 2018; Zhu et al., 2019) employed view interpolations (Chen and Williams, 1993) and depth-image-based rendering (Fehn, 2004) , but these suffer from interpolation artifacts, and pixel misalignment due to the use of noisy depth maps. Recent methods like Quark (Flynn et al., 2024) and SIMPLI (Solovev et al., 2023) employ a dynamic layered depth map (LDM) of the scene estimated through a deep neural network to perform view synthesis. However, these methods still require multi-view input, resulting in requirement of multiple capture devices and high communication bandwidth. Therefore, generating new views from a single input image and a specified camera pose has emerged as a key task in rendering. A variety of approaches, diffusion based 3DiM (Watson et al., 2022) and transformer-based architectures like NViST (Jang and Agapito, 2024), are proposed to tackle this problem. These methods leverage large scale datasets to be trained and produce high-quality novel views. While these approaches achieve remarkable fidelity, they are either unsuitable for real-time deployment or assume specific scene representations (e.g., dense layers or precomputed geometry). Our work builds on this trajectory by focusing on generalizable, feed forward synthesis that supports both real-time constraints and 6DoF pose inputs.
Real-time rendering is critical for telepresence and AR applications, but most state-of-the-art methods either depend on heavy iterative models (e.g., diffusion (Watson et al., 2022; Deitke et al., 2023)) or rely on attention-heavy transformer blocks (Jang and Agapito, 2024), both of which impede latency-sensitive deployments.
In this work, we build upon the growing body of research to address the challenges of computation efficiency and low latency for view synthesis. We propose a method designed to generate high-quality novel views from a single input image, named PVSNet, targeting real-time performance. PVSNet utilizes a position-aware rendering approach that efficiently handles camera movements in the scene, enabling smooth and realistic motion parallax for depth perception without the computational burden of traditional methods. In summary, our main contributions are as follows:
A neural network for real-time view synthesis from a single image and a target pose, supporting both translation only and full 6DoF inputs, making it suitable for telepresence applications.
@@
A compact and flexible positional embedding module similar to the one used in NeRF (Mildenhall et al., 2021), which is combined with a multi-layer perceptron (MLP), to create a more generalizable framework capable of handling different scenes.
An in-depth evaluation of the proposed method, demonstrating real-time performance with inference times of over 50 FPS at 512512 resolution, significantly outperforming existing single view methods in speed and image quality.
@@
The remainder of this paper is organized as follows: Section 2 reviews the state-of-the-art methods aimed at view synthesis. Section 3 describes the proposed approach for view synthesis. Section 4 provides a comparative analysis of the results, including quality and speed evaluations. Section 5 presents a detailed ablation study to analyze different components of the proposed method. Finally, conclusions are summarized in Section 6.
@@ 2. Background
View synthesis has recently become a pivotal research area in photorealistic immersive telepresence applications, aiming to generate unseen perspectives of a remote scene from a limited set of input images. Traditional approaches, such as structure-from-motion (SfM) and multi-view stereo (MVS), rely on geometric consistency between input views to reconstruct novel viewpoints (Seitz and Dyer, 1999). While these methods have shown effectiveness in controlled environments, their performance deteriorates in complex real-world scenarios and often necessitates multiple input views to achieve high-quality results. In light of these limitations, more recent methods focus on synthesizing novel views from sparse or single input images. This section outlines these advancements and their respective contributions to the field of view synthesis based on the number of input images required to perform view synthesis.
2.1. Multi-View Input
@@
Multi-view input based methods rely on several images or views of a scene to generate novel viewpoints. Recent advancements in multi-view input based methods have centered around leveraging neural implicit representations such as Neural Radiance Fields (NeRF) (Mildenhall et al., 2021). A NeRF encodes a scene as a continuous volumetric function, allowing novel view generation through volumetric rendering. Despite its capacity to produce highly photorealistic images, NeRF suffers from computational inefficiencies, particularly in terms of inference time. Several techniques have been proposed to mitigate these limitations, such as hash encoding in InstantNGP (Müller et al., 2022) and multi-scale representations in Mip-NeRF (Barron et al., 2021). While these improvements accelerate rendering, they retain significant reliance on MLPs, which remain computationally expensive.
@@
Plenoxels (Fridovich-Keil et al., 2022), which leverage spherical harmonics to represent a scene without the need for deep neural networks, have shown promising results when considering their fast optimization time. Though Plenoxels achieve faster optimization than many NeRF-based methods, it still falls short in terms of memory efficiency. A more recent innovation, Gaussian Splatting (Kerbl et al., 2023), improves rendering efficiency by employing high-dimensional Gaussian kernels to represent the scene. However, its optimal performance still depends on the availability of multiple input views, and the scenes with high frequency details require more kernels, increasing overall memory requirement.
@@
Light Field Rendering emerges as an alternative of scene regression methods like NeRF (Mildenhall et al., 2021), Plenoxels (Fridovich-Keil et al., 2022), and Gaussian Splatting (Kerbl et al., 2023). Light field (LF) representation encodes both the spatial and angular information of a scene, often described by an HxWxNxM structure, where HxW refers to the spatial resolution and NxM denotes the angular resolution (Levoy and Hanrahan, 1996). This representation facilitates tasks like viewpoint manipulation (Zhou et al., 2020). LF reconstruction methods aim to generate a dense NxM light field from few input views, eliminating the need for complex multi-camera setups. Common LF reconstruction methods (Zhou et al., 2020, 2021) utilize convolutional neural networks (CNNs) to estimate scene depth followed by a warping operator to reconstruct the full LF. To generate the novel views from the reconstructed LFs, algorithms like dynamically reparameterized LFs (Isaksen et al., 2000) and per-view disk-based blending (Overbeck et al., 2018) have to be employed. However, these methods often have a higher memory footprint, making the process unsuitable for dynamic environments or systems with limited memory.
Despite the effectiveness of LF-based methods, they also face challenges when working in live applications, where real-time performance and memory efficiency are critical. Recent works have introduced more efficient solutions to overcome these issues, such as SIMPLI (Solovev et al., 2023), GNPR (Suhail et al., 2022), and Quark (Flynn et al., 2024). SIMPLI utilizes a self-improving multiplane-to-layer image approach, enhancing view synthesis by leveraging a deep refinement network for scene representation. However, it still faces challenges in achieving fast inference speeds while maintaining high-quality results. GNPR, on the other hand, uses a generalizable neural framework that extracts patches from neighboring images along the epipolar lines to generate novel views. This method has demonstrated impressive quality, but like SIMPLI, it struggles with high computational demands and inefficiencies during inference. In contrast, Quark introduces a novel layered depth map (LDM) approach that combines efficiency and quality by performing a multi-scale iterative render-and-refine process. This method achieves real-time, high-resolution rendering, surpassing SIMPLI and GNPR in terms of both speed and quality. Despite its impressive performance, Quark remains limited to multi-view inputs, restricting its applicability in scenarios where only a single input image is available.
Our proposed method, PVSNet described in Section 3, goes a step further by offering a more efficient view synthesis approach, utilizing only a single input image. By removing the dependency on multiple views, PVSNet not only reduces the computational overhead, but also extends the applicability of view synthesis methods to real-time environments. In addition, the method ensures that high-quality novel views are generated while maintaining efficiency, making it suitable for resource-constrained applications. Unlike methods that rely on fixed view grids or dense multi-camera capture, PVSNet operates on arbitrary camera positions and orientations, making it suitable for unconstrained, real-time scenarios. The ability to embed full 6DoF poses directly into the synthesis pipeline allows our method to generalize across both panoramic and planar input modalities without requiring view-specific encodings or ray models.
2.2. Single-View Input
Single-view input methods aim to synthesize novel views from just one image, primarily addressing the scarcity of limited input data. Many of these techniques employ deep learning models to infer depth or scene structure from a single image. For instance, PixelNeRF (Yu et al., 2021) and VisionNeRF (Lin et al., 2023) generalize NeRF to new scenes by conditioning on image features at large scale. Although the PixelNeRF and VisionNeRF show promising results, they face challenges such as scale ambiguity and large baseline camera movements, which degrade rendering quality in more extreme viewpoints.
@@ 2.2. Single-View Input
To mitigate issues like scale ambiguity and substantial viewpoint changes, diffusion-based approaches have emerged. For example, 3DiM (Watson et al., 2022) employs pose guided diffusion to improve performance, particularly in scenarios with large translational movement. Fine-tuning diffusion models on large-scale synthetic datasets (Deitke et al., 2023) has also shown potential (Liu et al., 2023, 2024) in handling the scale ambiguity, although these methods are hindered by slow inference times, and highest memory footprint compared to all other methods discussed so far.
Other single-view input methods such as 3D Photo (Shih et al., 2020) and SLIDE (Jampani et al., 2021) focus on inferring views through learned layered depth inpainting. These methods can yield visually compelling results, especially for large baselines, but often require RGB+D input (Shih et al., 2020) or rely on multi-stage pipelines with monocular depth estimation, layering, and hole-filling (Jampani et al., 2021), making them unsuitable for real-time applications.
@@
As an alternative to diffusion-based frameworks, multiplane image (MPI) representations approximate a scene using a series of parallel planes at fixed depths, each containing a color and alpha map. This representation allows view synthesis based on a fixed (Zhou et al., 2018; Flynn et al., 2019; Tucker and Snavely, 2020; Li and Kalantari, 2020) or variable (Li and Kalantari, 2020) number of planes. While MPI-based methods achieve superior visual quality, their computational demands remain high, especially during MPI layer prediction.
More recent innovations, such as NViST (Jang and Agapito, 2024), incorporate vision transformers (ViTs) (Dosovitskiy, 2020) into view synthesis architectures, demonstrating strong generalization capabilities in real-world settings. By leveraging masked autoencoders (MAEs) and adaptive layer normalization conditioned on camera parameters, NViST delivers high-quality results, even for out-of-distribution views. Despite these strengths, the reliance of NViST on a ViT backend and patch-wise image processing leads to slower inference at higher resolutions. Meanwhile, large-scale GANs (Koh et al., 2023) offer rapid rendering and handle large baselines effectively at high resolution from a single ODI, but it also incurs significant memory overhead, limiting the practical usage in resource-constrained settings. Another relevant line of work is MegaParallax (Bertel et al., 2019), which aims to produce motion parallax in ODI using layered mesh generated from monocular depth maps and image warping. While it succeeds in synthesizing novel viewpoints from a single ODI, the pipeline involves explicit depth estimation, foreground segmentation, and layered rendering, which introduce computational overhead and make real-time operation impossible.
In summary, despite the advances, single-view methods frequently grapple with achieving both real-time performance and high-quality novel view generation in live applications. Addressing this balance between efficiency and fidelity remains a critical goal for immersive telepresence and related applications. The limitations identified here, which includes scale ambiguity, heavy computational demands, and memory constraints, inform the design of our proposed approach PVSNet (described in Section 3), which aims to reconcile the trade-offs these methods face.
3. Method: Position-Aware View-Synthesis
PVSNet is a two stage model which, given the source image and the target position , aims to synthesize the novel image at the corresponding target viewpoint. Our approach consists of two key stages as shown in Figure 2. First, the target position is embedded into a high-dimensional space through positional embedding which is discussed in Section 3.1, and second, new views are rendered from these positions, which is detailed in Section 3.2.